Suppose that a family has 5 children. Also, suppose that the probability of having girl is 1/2. Find the probabilities that the family has the following children,
a) no more than 4 girls

Answer by stanbon(48568)

Suppose that a family has 5 children.
Find the probabilities that the family has the following children,
a) no more than 4 girls
---------------
This is a binomial problem with n=5; p = 1/2; 0 <= x < 5
P(0 <= x <5) = 1 - P(x=5) = 1 - (1/2)^5 = 31/32
------------------
Cheers,
Stan H.